the square root of x-1 +x=3
Answer by checkley77(12844)

sqrt(x-1)+x=3
sqrt(x-1)+x=3
sqrt(x-1)=3-x square both sides.
x-1=9-6x+x^2
x^2-6x+9-x+1=0
x^2-7x+10=0
(x-5)(x-2)=0
x-5=0
x=5 ans.
x-2=0
x=2 ans.
Proof:
sqrt(5-1)+5=3
sqrt4=3-5
-2=-2
sqrt(2-1)+2=3
sqrt1=3-2
1=1